# UFT: Unifying Fine-Tuning of SFT and RLHF/DPO/UNA through a Generalized Implicit Reward Function

### v1

@@ Abstract
By pretraining on trillions of tokens, an LLM gains the capability of text generation. However, to enhance its utility and reduce potential harm, SFT and alignment are applied sequentially to the pretrained model. Due to the differing nature and objective functions of SFT and alignment, catastrophic forgetting has become a significant issue. To address this, we introduce Unified Fine-Tuning (UFT), which integrates SFT and alignment into a single training stage using the same objective and loss functions through an implicit reward function.
Our experimental results demonstrate that UFT outperforms SFT on instruction-tuning data alone. Moreover, when combining instruction-tuning data with alignment data, UFT effectively prevents catastrophic forgetting across these two stages and shows a clear advantage over sequentially applying SFT and alignment. This is evident in the significant improvements observed in the ifeval task for instruction-following and the truthful-qa task for factuality. The proposed general fine-tuning framework UFT establishes an effective and efficient pretraining-UFT paradigm for LLM training.
@@ 1 Introduction
For RLHF/PPO, a dataset is created consisting of triplets: a prompt, a desired response, and an undesired response. This preference dataset is then used to train a pointwise reward model that evaluates a pair of prompt and response using the Bradley-Terry (BT) model [BT52]. During the RL stage of RLHF, a multi-objective optimization is performed to balance maximizing the reward score from the pretrained reward model for a prompt-response pair and minimizing the divergence between the fine-tuned model and the original pretrained model using KL divergence. However, RLHF is known for being unstable and computationally expensive. To address these issues, DPO proposes a method to optimize the reward model and the policy model together, transforming the original two-stage task in RLHF into a single-stage classification process. Although DPO simplifies the training task of RLHF, it only utilizes responses as either desired or undesired, lacking the more nuanced reward scores provided by reward models. KTO extends DPO slightly by replacing preference feedback with binary feedback. Recognizing the limitations that DPO can only handle pairwise feedback and KTO can only handle binary feedback, and they are both significantly different from RLHF/PPO, UNA proposes a unified framework that combines these alignment methods and various types of feedback through a generalized implicit reward function, transforming the unstable RLHF into a stable supervised learning problem.
For pretrained LLMs, SFT and alignment are typically applied in sequential order. This sequential approach often leads to catastrophic forgetting, where the model loses capabilities acquired in earlier stages. To address this, ORPO introduced a new objective, but it is constrained by its reliance on pairwise feedback and its limitation on effectiveness [HLT24]. On the other hand, PAFT proposed applying SFT and alignment in parallel, followed by merging these adaptors with the pretrained model [PWB+24]. However, the PAFT method is cumbersome as it involves three stages: SFT, alignment, and model merging and it requires sparsity of SFT and alignment to avoid interference during merging. This paper aims to tackle the catastrophic forgetting problem by introducing a mathematically proven and efficient methodology, i.e., UFT.
We are inspired by UNA’s ability to process score-based feedback effectively. Consequently, we aim to extend UNA to achieve the objectives of SFT. Specifically, SFT involves maximizing the probability of a response given an instruction. Therefore, instruction-tuning data can be considered alignment data with the highest feedback score, such as the desired response for pairwise feedback, a thumbs-up for binary feedback, and a score of 1 for score-based feedback. In the paper, we have demonstrated that UFT and SFT both aim to maximize the likelihood of the responses in instruction-tuning data. By combining the instruction-tuning dataset with the alignment data, we can fine-tune a pretrained LLM with the same objective and loss function while effectively preventing catastrophic forgetting. Our experiments demonstrate that UNA, when applied to instruction-tuning data, can outperform traditional SFT on the same data in downstream tasks. Furthermore, this new perspective allows us to mix the instruction-tuning dataset with the alignment dataset, enabling us to fine-tune LLMs in a single step. Our experiments indicate that this approach prevents the issue of catastrophic forgetting by surpassesing the performance of the previous sequential training pipeline. This is evident in the significant improvements observed in the ifeval task for instruction-following [ZLM+23] and the truthful-qa task for factuality [LHE22].
Lastly, we have conducted a direct comparison between UFT and the pretraining stage. Drawing from an old Chinese proverb, ”Read ten thousand books, travel ten thousand miles”, we can liken the training of a LLM to a successful human life as shown in Figure. 1. The pretraining stage corresponds to ”Read ten thousand books,” where the model is trained on billions of tokens to predict the next token without receiving any feedback. In contrast, the fine-tuning stage is akin to ”Travel ten thousand miles,” where the LLM is posed with various questions, generates responses, and receives feedback from different labelers and utilize these feedback for improvement. The existing fine-tuning paradigm involves 2-stage training: SFT followed by alignment. Our new UFT unifies the two stages, leading to a post-training framework parallel to pretraining.
@@
    UFT that unifies SFT and Alignment solves the catastrophic forgetting problem elegantly and surpasses the performance of the original sequential application order in downstream tasks.
@@ 2.2 RLHF, DPO and UNA
The training of RLHF is memory-intensive because it requires maintaining both the explicit reward model and the policy model. Additionally, reinforcement learning is notorious for its instability. To address this issue, DPO proposes creating a mapping between the optimal policy and the reward model, i.e., an implicit reward model, as illustrated in Eq. 4, and optimizing them together. However, $Z(x)$ is intractable and can only be canceled out by subtracting the implicit reward of the desired response $y_{w}$, i.e., $r_{\theta}(x,y_{w})$ from the implicit reward of the undesired response $y_{l}$, i.e., $r_{\theta}(x,y_{l})$. This limitation confines DPO to pairwise datasets only. Furthermore, DPO cannot utilize the precise evaluation from the explicit reward model in RLHF. KTO extends DPO to binary feedback by estimating $Z(x)$ from multiple responses to the same prompt.
@@
To address this issue, UNA offers an alternative proof and demonstrates a novel mapping between the optimal policy and the reward model, i.e., the generalized implicit reward model. The form of the mapping between the generalized implicit reward model and policy is $r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)+f(x)+c$ where $f(x)$ measures the quality of the prompts and $c$ adjusts the offset between implicit and explicit rewards. With simplification, i.e., $f(x)=c=0$, the relationship between the generalized implicit reward model and policy can be shown in Eq. 5. Unlike DPO, which is constrained by $Z(x)$ to pairwise feedback, UNA is versatile enough to handle various types of data, including pairwise feedback, binary feedback, and score-based feedback. These data types are optimized by minimizing the difference between the implicit reward in Eq. 5 and the explicit reward $r_{\phi}(x,y)$, which is provided by human labelers, other LLMs, or the explicit reward model, as shown in Eq. 6. When LLMs and explicit reward models are used to evaluate responses in real-time, it is referred to as online UNA. Conversely, if feedback is labeled beforehand, it is termed offline UNA. Consequently, UNA can function in both online and offline modes, effectively bridging the gap between online RLHF and offline DPO and KTO.
@@
Figure 2: Subfigure (A) refers to SFT, Subfigure (B) refers to Alignment including RLHF, DPO and UNA and Subfigure (C) refers to UFT. Traditionally, the fine-tuning process begins with SFT followed by alignment. However, the proposed UFT method integrates both SFT and alignment into a single, cohesive process. Additionally, UFT can be adapted for online implementation, provided that the feedback scores are derived from reward models based on real-time prompts and responses, rather than relying on pre-collected feedback.
2.3 UFT Unifying SFT and Alignment
To enhance the capabilities of LLMs in question answering and addressing ethical issues, SFT and alignment are typically applied in a sequential manner due to the distinct nature of these tasks. However, this sequential approach often leads to the well-known issue of catastrophic forgetting. In this study, we introduce UFT which integrates SFT and alignment into a single stage, thereby mitigating this problem.
To be more specific, the instruction-tuning data consist of a prompt $x$ and a corresponding response $y$, where the response $y$ is considered of high quality, typically labeled by experts in the field. In comparison, the alignment data include a prompt $x$, response $y$, and feedback $r$. This feedback can be categorized as desired or undesired for pairwise feedback, positive or negative for binary feedback, or a scalar in the interval [0, 1] for general score-based feedback. Due to the high quality of instruction-tuning data, they can be regarded as data with a score of 1, i.e., positive feedback. With this consideration, the instruction-tuning dataset can be transformed into alignment data in the format of prompt $x$, response $y$ and feedback $r=1$, which is the highest reward in consideration. Eventually, the transformed instruction-tuning dataset and the alignment dataset can be merged for fine-tuning LLM using the loss function in UNA as shown in Eq. 6. Our experiments demonstrate that when fine-tuning using only instruction-tuning dataset, UFT can outperform SFT on downstream tasks, which we attribute to the KL divergence term that focuses on minimization with the pretrained model, a factor ignored in the SFT processes. Additionally, the results indicate that mixing the instruction-tuning data with alignment for UFT can prevent the catastrophic forgetting problem and outperform previous sequential methods. Lastly, we discover that the distribution of mixed data, i.e., the proportion of instruction-tuning data and alignment data will impact the performances of LLMs. More details can be found in the experiment section.
A non-rigorous proof why UFT can replace SFT is provided by arguing that they achieve the same goal of maximizing the probability of $\pi_{\theta}(y|x)$ for prompt $x$ and response $y$ in the instruction-tuning dataset. For SFT, the probability of $\pi_{\theta}(y|x)$ is directly maximized by cross entropy-loss. In UFT, suppose we apply the Sigmoid function on the implicit reward in Eq. 5 and using mean square error (MSE) to measure the difference of implicit reward and explicit reward, the objective will become Eq. 7. This loss function will drive $r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)$ to positive infinity. Since both $\beta$ and $\pi_{\text{ref}}(y|x)$ are fixed, this objective will maximize $\pi_{\theta}(y|x)$. As a result, UFT can replace SFT to maximize the probability of generating the response, and it outperforms SFT by minimizing the difference to the pretrained model.
3 Experiments
@@
Model gsm8k truthful-qa winograde arc hellaswag mmlu average
@@
3.1 UFT vs. SFT
To begin with, we compare UFT with SFT on the same instruction-tuning dataset. The UltraChat dataset is utilized by unfolding one conversation into multiple training examples to faciliate the training of UFT. From the UltraChat dataset, 20k samples are selected and utilized for training. The Mistral 7B-v0.1 [JSM+23] is utilized as the base model, with low rank adaptation (LoRA) of $r=16$ [HSW+21]. For SFT, the learning rate of $1e^{-4}$ is the best, and for UFT, the learning rate of $3e^{-5}$ and $\beta$ of 0.01 is the best.
The fine-tuned models are tested on different tasks including 12 tasks on two HuggingFace Open LLM Leaderboards[BFH+23, FHL+24]. The new Open LLM Leaderboard includes 6 tasks: bbh (Suzgun et al., 2022), gpqa (Rein et al., 2023), mmlu-pro (Wang et al., 2024), musr (Sprague et al., 2024), ifeval (Zhou et al., 2023), and math-hard (Hendrycks et al., 2021). For all these tasks, the average scores are reported. Conversely, the old Open LLM Leaderboard comprises 6 different tasks: gsm8k (Cobbe et al., 2021), truthful-qa (Lin et al., 2022), winograde (Sakaguchi et al., 2019), arc (Allen AI), hellaswag (Zellers et al., 2019), and mmlu (Hendrycks et al., 2021). In this work, the average match rate in gsm8k, mc2 in truthful-qa, accuracy in winograde, acc-norm in arc, acc-norm in hellaswag, and accuracy in mmlu will be reported. Additionally, MT-Bench [ZCS+23] and Alpaca-eval [LZD+23] will be used to evaluate the model’s ability to generate text responses, rather than selecting from predefined candidate answers. Specifically, the Length-Controlled Win Rate (LC WR) will be documented for Alpaca-eval. Additionally, the average length of the outputs within this evaluation will be provided for comprehensive analysis.
As demonstrated in Table 1 and Table 2, Mistral+UFT surpasses Mistral+SFT in 8 out of 12 tasks. When considering the average performances across these leaderboards, Mistral+UFT consistently outperforms Mistral+SFT. This indicates that UFT enhances the capabilities of an LLM by maximizing the reward and minimizing the difference from the pretrained model. Furthermore, in the MT-Bench evaluation, Mistral+UFT outperforms Mistral+SFT, whereas in the Alpaca-eval, Mistral+SFT overshadows Mistral+UFT. From a generation capability standpoint, both UFT and SFT exhibit similar performance. As a result, when evaluating the performance on downstream tasks, UFT demonstrates superiority over SFT.
3.2 Catastrophic Forgetting Study
Another advantage of UFT lies in combining SFT and alignment into one stage to avoid catastrophic forgetting. In this stage, we utilize the HelpSteer2 dataset of 20k examples for alignment [WDD+24]. For UFT, the 20k examples from UltraChat and 20k examples from HelpSteer2 are merged and utilized for training. For comparison, the best performing SFT model of learning rate $1e^{-4}$ in the previous experiment is utilized for further fine-tuning using DPO, KTO and UNA.
The same tasks are utilized for evaluation. In the two HuggingFace open LLM leaderboards, Mistral+UFT outperforms all of Mistral+SFT+DPO, Mistral+SFT+KTO, and Mistral+SFT+UNA in 9 out of 12 tasks. In terms of average scores, Mistral+UFT surpasses all three sequential methods. Several aspects need further discussion. Firstly, the catastrophic forgetting problem is evident as the performances of Mistral+SFT+DPO, Mistral+SFT+KTO, and Mistral+SFT+UNA are worse than Mistral+SFT, a phenomenon also known as alignment tax. Additionally, we observe that Mistral+UFT shows significant improvements on ifeval, which tests the model’s capability of instruction-following, and truthful-qa, which assesses the model’s alignment capabilities. These results indicate that UFT greatly enhances instruction-following and alignment capabilities, demonstrating its effectiveness. In terms of generation capability, Mistral+UFT outperforms the other three sequential methods and it does not seem to suffer from catastrophic forgetting, as it performs better than both Mistral+SFT and Mistral+UFT on instruction-tuning data alone. Moreover, Mistral+UFT does not bias towards long generation like the other three sequential methods, which is another advantage of UFT.
3.3 Data Distribution in Training UFT
Several intriguing points will be discussed. As shown in Table 4 and Table 5, among the 12 tasks, one task, namely musr, shows a minor statistical decrease in performance. In contrast, four tasks-gpqa, gsm8k, winograde, and arc—exhibit statistically small improvements. Notably, two tasks, ifeval and truthful-qa, demonstrate statistically significant improvements. Specifically, ifeval improves from 23.22 to around 44, and truthful-qa improves from 42.58 to around 53. Significant advancements are observed in MT-Bench and Alpaca-eval, indicating a substantial enhancement in the generation capability of the LLM as shown in Table 6. This aligns with our expectations, as instruction-tuning data primarily aim to enhance the model’s instruction-following capabilities, which positively impacts ifeval, MT-Bench, and Alpaca-eval. Meanwhile, alignment data contribute to improvements in tasks like truthful-qa. Other tasks remain largely unaffected due to the absence of relevant data, suggesting that incorporating more pertinent data could enhance their performance. When increasing the proportion of instruction-tuningg data, we observe a decrease in truthful-qa performance from 56 to 50, indicating that a larger proportion of alignment data benefits the bias and ethical performance of the LLM. On the other hand, simply adding more instruction-tuning data does not improve the performance of ifeval, MT-Bench, and Alpaca-eval. Therefore, further investigation into the optimal ratio between instruction-tuning data and alignment data is warranted.

### v2

@@ Abstract
By pretraining on trillions of tokens, an LLM gains the capability of text generation. However, to enhance its utility and reduce potential harm, SFT and alignment are applied sequentially to the pretrained model. Because SFT and alignment have different objectives and underlying processes, performance on certain tasks can decline. To address this, we introduce Unified Fine-Tuning (UFT), which integrates SFT and alignment into a single training stage using the same objective and loss functions through an implicit reward function.
Our experimental results demonstrate that UFT outperforms SFT on instruction-tuning data alone. Moreover, when combining instruction-tuning data with alignment data, UFT effectively prevents the degradation on some tasks across these two stages and shows a clear advantage over sequentially applying SFT and alignment. This is evident in the significant improvements observed in the ifeval task for instruction-following and the truthful task for factuality. The proposed general fine-tuning framework UFT establishes an effective and efficient pretraining-UFT paradigm for LLM training.
@@ 1 Introduction
For pretrained LLMs, SFT and alignment are traditionally performed in sequence. However, this staged approach often leads to performance degradation, where the model loses capabilities acquired in earlier phases. This paper seeks to address and mitigate this degradation.
Drawing inspiration from UNA’s effectiveness in handling score-based feedback, we propose extending UNA to also fulfill the goals of SFT. By converting SFT data into a format compatible with alignment training, we enable the use of a unified objective and loss function. This approach allows for effective fine-tuning of pretrained LLMs while preserving their previously acquired capabilities across tasks.
@@
    UFT that unifies SFT and alignment solves the performance degradation on some tasks elegantly and surpasses the performance of the original sequential application order in downstream tasks.
@@ 2.2 RLHF, DPO and UNA
The training of RLHF is memory-intensive because it requires maintaining both the explicit reward model, the value model and the policy model. Additionally, reinforcement learning is notorious for its instability.
To address this issue, DPO proposes creating a mapping between the optimal policy and the reward model, i.e., an implicit reward model, as illustrated in Eq. 4, and optimizing them together. However, $Z(x)$ is intractable and can only be canceled out by subtracting the implicit reward of the desired response $y_{w}$, i.e., $r_{\theta}(x,y_{w})$ from the implicit reward of the undesired response $y_{l}$, i.e., $r_{\theta}(x,y_{l})$. This limitation confines DPO to pairwise datasets only. Furthermore, DPO cannot utilize the precise evaluation from the explicit reward model in RLHF. KTO extends DPO to binary feedback by estimating $Z(x)$ from multiple responses to the same prompt.
@@
To address this issue, UNA offers an alternative proof and demonstrates a novel mapping between the optimal policy and the reward model, i.e., the generalized implicit reward model. The form of the mapping between the generalized implicit reward model and policy is shown in Eq. 5. Unlike DPO, which is constrained by $Z(x)$ to pairwise feedback, UNA is versatile enough to handle various types of data, including pairwise feedback, binary feedback, and score-based feedback. These data types are optimized by minimizing the difference between the implicit reward in Eq. 5 and the explicit reward $r_{\phi}(x,y)$, which is provided by human labelers, other LLMs, or the explicit reward model, as shown in Eq. 6. When LLMs and explicit reward models are used to evaluate responses in real-time, it is referred to as online UNA. Conversely, if feedback is labeled beforehand, it is termed offline UNA. Consequently, UNA can function in both online and offline modes, effectively bridging the gap between online RLHF and offline DPO and KTO.
@@
Figure 2: Subfigure (A) refers to SFT, Subfigure (B) refers to alignment including RLHF, DPO and UNA and Subfigure (C) refers to UFT. Traditionally, the fine-tuning process begins with SFT followed by alignment. However, the proposed UFT method integrates both SFT and alignment into a single, cohesive process.
@@
Model gsm8k truthful winograde arc hellaswag mmlu average
@@
2.3 UFT: Unify SFT and Alignment
To enhance the capabilities of LLMs in question answering and addressing ethical issues, SFT and alignment are typically applied in a sequential manner due to the distinct nature of these tasks. However, this sequential approach often leads to performance degradation on some tasks. In this study, we introduce UFT which integrates SFT and alignment into a single stage, thereby mitigating this problem.
To be more specific, the instruction-tuning data consist of a prompt $x$ and a corresponding response $y$, where the response $y$ is considered of high quality, typically labeled by experts in the field. In comparison, the alignment data include a prompt $x$, response $y$, and feedback $r$. This feedback can be categorized as desired or undesired for pairwise feedback, positive or negative for binary feedback, or a scalar in the interval [0, 1] for general score-based feedback. Due to the high quality of instruction-tuning data, they can be regarded as data with a score of 1, i.e., positive feedback. With this consideration, the instruction-tuning dataset can be transformed into alignment data in the format of prompt $x$, response $y$ and feedback $r=1$, which is the highest reward in consideration. Eventually, the transformed instruction-tuning dataset and the alignment dataset can be merged for fine-tuning LLM using the loss function in UNA as shown in Eq. 6. Our experiments demonstrate that when fine-tuning using only instruction-tuning dataset, UFT can outperform SFT on downstream tasks, which we attribute to the KL divergence term that focuses on minimization with the pretrained model, a factor ignored in the SFT processes. Additionally, the results indicate that mixing the instruction-tuning data with alignment for UFT can prevent the performance degradation on some tasks and outperform previous sequential methods. Lastly, we discover that the distribution of mixed data, i.e., the proportion of instruction-tuning data and alignment data will impact the performances of LLMs. More details can be found in the experiment section.
A heuristic proof why UFT can replace SFT is provided by arguing that they achieve the same goal of maximizing the probability of $\pi_{\theta}(y|x)$ for prompt $x$ and response $y$ in the instruction-tuning dataset. For SFT, the probability of $\pi_{\theta}(y|x)$ is directly maximized by cross entropy-loss. In UFT, suppose we apply the Sigmoid function on the implicit reward in Eq. 5 and using mean square error (MSE) to measure the difference of implicit reward and explicit reward, the objective will become Eq. 7. This loss function will drive $r_{\theta}(x,y)=\beta\log\left(\frac{\pi_{\theta}(y|x)}{\pi_{\text{ref}}(y|x)}\right)$ to positive infinity. Since both $\beta$ and $\pi_{\text{ref}}(y|x)$ are fixed, this objective will maximize $\pi_{\theta}(y|x)$. As a result, UFT can replace SFT to maximize the probability of generating the response, and it outperforms SFT by minimizing the difference to the pretrained model.
2.4 Comparison of Pretraining and Fine-Tuning
By integrating SFT and alignment through UFT, we can establish a unified fine-tuning framework that runs parallel to the pretraining phase. Drawing inspiration from the old Chinese proverb ”Read ten thousand books, travel ten thousand miles”, we can liken the pretraining stage to ”Read ten thousand books”. In this phase, a LLM absorbs and learns from an extensive collection of tokens, amounting to trillions. The primary objective during this stage is to predict the next token based on the previous context window, optimizing the model’s performance through the cross entropy loss function. Conversely, the fine-tuning stage can be likened to the adage ”Travel ten thousand miles,” where the model is exposed to a wide array of prompts. It generates diverse responses and learns from the feedback it receives, thereby refining and enhancing its capabilities.
Model bbh gpqa mmlu-pro musr ifeval math-hard average
Qwen 67.89 37.32 57.93 49.12 42.98 31.63 47.81
Qwen+SFT 67.75 40.42 58.13 48.73 47.48 30.77 48.88
Qwen+UFT 67.67 40.42 58.1 48.47 46.36 31.23 48.71
Qwen+SFT+DPO 67.8 39.01 58.06 49.4 47.48 33.67 49.24
Mistral+SFT+KTO 67.61 39.26 58.19 48.34 53.03 35.07 50.27
Mistral+SFT+UNA 68.84 38.42 58.35 49.14 57.33 34.7 51.13
Mistral+UFT 67.73 38.59 58.73 47.42 64.05 37.8 52.39
Table 4: Comparison of Qwen+SFT and Qwen+UFT on instruction-tuning data, and comparison of Qwen+SFT+DPO, Qwen+SFT+KTO, Qwen+SFT+UNA, and Qwen+UFT on both instruction-tuning and alignment data on the new HuggingFace open LLM Leaderboard
Model gsm8k truthful winograde arc hellaswag mmlu average
Qwen 85.79 57.75 82 70.31 85.22 83.19 77.38
Qwen+SFT 87.38 58.81 82 70.9 84.99 83.39 77.91
Qwen+UFT 88.29 58.96 82.64 70.9 85 83.34 78.19
Qwen+SFT+DPO 85.67 58.54 82 70.14 85.24 83.31 77.48
Qwen+SFT+KTO 85.56 59.77 82.16 70.14 85.32 83.32 77.71
Qwen+SFT+UNA 92.08 62.74 82.4 71.5 85.37 83.36 79.58
Qwen+UFT 90.68 66.7 83.58 71.84 85.58 83.35 80.29
Table 5: Comparison of Qwen+SFT and Qwen+UFT on instruction-tuning data, and comparison of Qwen+SFT+DPO, Qwen+SFT+KTO, Qwen+SFT+UNA, and Qwen+UFT on both instruction-tuning and alignment data on the old HuggingFace open LLM Leaderboard
Model MT-Bench Alpaca-eval
LC WR Length
Qwen 7.97 8.47 2176
Qwen+SFT 7.85 8.34 2800
Qwen+UFT 7.95 9.96 1292
Qwen+SFT+DPO 8.64 9.83 2664
Qwen+SFT+KTO 8.58 10.17 2149
Qwen+SFT+UNA 8.57 13.75 1328
Qwen+UFT 8.67 13.79 1307
Table 6: Comparison of Qwen+SFT and Qwen+UFT on instruction-tuning data, and comparison of Qwen+SFT+DPO, Qwen+SFT+KTO, Qwen+SFT+UNA, and Qwen+UFT on both instruction-tuning and alignment data using MT-Bench and Alpaca-eval
3 Experiments